OpenStudy (anonymous):

Kean read 3 books in 4 weeks. How long would it take him to read 12 books?

OpenStudy (cwrw238):

have you missed out the number of books in your second sentence?

OpenStudy (shubhamsrg):

lol..

OpenStudy (anonymous):

o 1 minute

OpenStudy (shubhamsrg):

well with that speed,, she'll read books in days.. as / = / and we apply unitary method!! ;)

OpenStudy (anonymous):

3 book = 4 weeks 1 book = 4/3 weeks 12 books = 4/3 * 12 = 16 weeks